0 0
Add a comment Improve this question Transcribed image text
Answer #1

if equale sides are a,

W = Σ(k*q1*q2/r)

W = k*Σ(q1*q2/r)

= 9e9*((23.8e-6*23.8e-6/2) + (23.8e-6*79e-6/2) + (23.8e-6*79e-6/2))

W = 19.47 J

however, if equal sides are b,

W = Σ(k*q1*q2/r)

W = k*Σ(q1*q2/r)

= 9e9*((23.8e-6*23.8e-6/2) + (23.8e-6*79e-6/2) + (23.8e-6*79e-6/2))

W = 19.47 J